OpenStudy (anonymous):

Can someone help me solve this sqrt 50x^7y^7 * sqrt 6xy^4

OpenStudy (anonymous):

\[\sqrt 50x^7y^7 * \sqrt 6xy^4\] \[\sqrt{50*6} x^{7+1} y^{7+4}\]

OpenStudy (shamim):

is it \[\sqrt{50x ^{7}y ^{7}}\sqrt{6xy ^{4}}\]

OpenStudy (anonymous):

sqrt 50x^7y^7 * sqrt 6xy^4

OpenStudy (shamim):

\[or, \sqrt{50 \times 6 x ^{8}y ^{11}}\]

OpenStudy (shamim):

\[=25\sqrt{3}x ^{4}y ^{5}\sqrt{y}\]

OpenStudy (shamim):

this is ur result

OpenStudy (anonymous):

I'm lost

OpenStudy (shamim):

did u not get it

OpenStudy (anonymous):

How do you show the steps

OpenStudy (shamim):

ok i m telling again

OpenStudy (anonymous):

sqrt{50*6} x^{7+1} y^{7+4} what is this someone else put this up

OpenStudy (shamim):

\[\sqrt{50x ^{7}y ^{7}}\sqrt{6xy ^{4}}=\sqrt{50x ^{7} y ^{7} \times 6xy ^{4}}=\sqrt{50 \times 6x ^{8}y ^{11}}\]

OpenStudy (shamim):

ya u r right

OpenStudy (shamim):

should i proceed further

OpenStudy (anonymous):

yes so that would be the first step you did or what I posted

OpenStudy (shamim):

ya i started frm ur posted question

OpenStudy (shamim):

\[\sqrt{2 \times 3 \times 2 \times 5 \times 5y \times x ^{4} \times x ^{4} \times y ^{5} \times y ^{5}}\]

OpenStudy (anonymous):

now im definitely lost just when i thought i was understanding

OpenStudy (shamim):

\[=2 \times 5x ^{4}y ^{5} \sqrt{3 \times y}\]

OpenStudy (shamim):

anyway which country u r frm

OpenStudy (shamim):

may i know ur country name. just for my curisity. nothing more

OpenStudy (anonymous):

this is what Ive got so far sqrt 50*6 x^(7+1)y^(7+4) sqrt 50*6 x^(8)y^(11) sqrt 2 * 3* 2* 5y * x^4 * x^4 * y^5 * y^5 sqrt 2 * 5x^4y^5 * sqrt 3 * y

OpenStudy (shamim):

its result. i think

OpenStudy (shamim):

i think the last step is our result. is not it

OpenStudy (shamim):

u r most welcome